ABC Company has set the following standards in order to produce one unit of its single product: standard quantity standard price direct materials 2.7 yards ?? per yard direct labor 2 hours $14 per hour variable overhead 2 hours ?? per hour During August, ABC Company spent $133,770 to purchase direct materials and had direct labor totaling $189,440. During August, ABC Company used 16,000 yards of direct materials in the production of units. ABC had 2,500 yards of materials in its August 1 direct material inventory and had 1,200 yards yards remaining in its direct materials inventory at August 31. ABC Company reported the following variances for August: Direct material price variance .............. $12,495 unfavorable Direct labor rate variance .................. $24,240 unfavorable Total direct labor variance ................. $6,560 favorable Variable overhead spending variance ......... $1,900 unfavorable Variable overhead efficiency variance ....... $20,900 favorable Calculate ABC Company's actual variable overhead cost for August.
Answer #1

Firstly we need to calculate actual direct labor hours and standard variable overhead cost which is shown as follows:-

Direct labor rate variance = (Std labor rate*Actual labor hours) - Actual direct labor cost

$24,240 Unfavorable = (Std labor rate*Actual labor hours) - $189,440

-24,240+189,440 = $14*Actual labor hours

14*Actual labor hours = $165,200

Actual labor hours = $165,200/14 = 11,800 hours

Total Direct Labor variance = Std Labor cost - Actual labor cost

$6,560 Favorable = Std Labor cost - $189,440

Std labor cost = 6,560+189,440 = $196,000

Std labor hours = Std labor cost/Std labor rate per hour

= $196,000/$14 per hour = 14,000 hours

Calculation of actual variable overhead rate per hour.

Variable Over.head efficiency variance = (Std hours - Actual hours)*Std variable overhead rate

20,900 favorable = (14,000 hrs - 11,800 hrs)*Std variable overhead rate

Std variable overhead rate = 20,900/(14,000 - 11,800) hrs

= 20,900/2,200 hrs = $9.50 per hour

Variable Overhead spending variance = (Std rate - Actual rate)*Actual hours

1,900 Unfavorable = ($9.50 - Actual rate)*11,800 hrs

-1,900/11,800 hrs = $9.50 - Actual rate

-0.161016949 = $9.50 - Actual rate

Actual rate = $9.50+0.161016949 = $9.661016949

Actual variable overhead cost = Actual labor hours*Actual variable overhead rate

= 11,800 hrs*$9.661016949 = $114,000

Therefore actual variable overhead cost for August is $114,000.